# Open the Eyes of MPNN: Vision Enhances MPNN in Link Prediction

### v1

@@ Abstract
Message-passing graph neural networks (MPNNs) and structural features (SFs) are cornerstones for the link prediction task. However, as a common and intuitive mode of understanding, the potential of visual perception has been overlooked in the MPNN community. For the first time, we equip MPNNs with vision structural awareness by proposing an effective framework called Graph Vision Network (GVN), along with a more efficient variant (E-GVN). Extensive empirical results demonstrate that with the proposed frameworks, GVN consistently benefits from the vision enhancement across seven link prediction datasets, including challenging large-scale graphs. Such improvements are compatible with existing state-of-the-art (SOTA) methods and GVNs achieve new SOTA results, thereby underscoring a promising novel direction for link prediction.
@@ 1 Introduction
Link prediction is a core task in graph machine learning that focuses on forecasting the presence of a connection between two nodes. This task is crucial in numerous applications such as recommendation systems (He et al., 2020), drug interaction prediction (Yamanishi et al., 2008), and knowledge-based reasoning (Bordes et al., 2013).
@@ 4.1 Scope-decoupled Subgraph Visualization
Decoupled Vision Scope. While the perception scope expands during message passing in MPNN, we fix the visual perception scope $k$ during visualization. This approach ensures that the scope in vision can be decoupled from which in message passing and offers two key advantages. First, both theoretical analysis and empirical evidence (Zeng et al., 2021) show that a decoupled scope independent from message-passing can improve expressivity of MPNN by addressing the over-smoothing issue (Li et al., 2018; Kenta Oono, 2020; Shen et al., 2024), and enhances the computational scalability of the models by alleviating the neighbor explosion problem (Chiang et al., 2019; Zeng et al., 2019).
@@
Second, as shown in Figure 2, an excessively large visual perception scope can lead to overcrowding of nodes/edges on a fixed-size canvas, resulting in a cluttered appearance that complicates graph interpretation. By using a small $k$, we maintain image clarity independent of message passing.

### v2

@@ Abstract
Message-passing graph neural networks (MPNNs) and structural features (SFs) are cornerstones for the link prediction task. However, as a common and intuitive mode of understanding, the potential of visual perception has been overlooked in the MPNN community. For the first time, we equip MPNNs with vision structural awareness by proposing an effective framework called Graph Vision Network (GVN), along with a more efficient variant (E-GVN). Extensive empirical results demonstrate that with the proposed frameworks, GVN consistently benefits from the vision enhancement across seven link prediction datasets, including challenging large-scale graphs. Such improvements are compatible with existing state-of-the-art (SOTA) methods and GVNs achieve new SOTA results, thereby underscoring a promising novel direction for link prediction. The official code is available at https://github.com/WEIYanbin1999/EGVN.
@@ 1 Introduction
Link prediction is a core task in graph machine learning that focuses on forecasting the presence of a connection between two nodes. This task is crucial in numerous applications such as recommendation systems (He et al., 2020), drug interaction prediction (Yamanishi et al., 2008), and knowledge-based reasoning (Bordes et al., 2013; Wei et al., 2023).
@@ 4.1 Scope-decoupled Subgraph Visualization
Decoupled Vision Scope. While the perception scope expands during message passing in MPNN, we fix the visual perception scope $k$ during visualization. This approach ensures that the scope in vision can be decoupled from which in message passing and offers two key advantages. First, both theoretical analysis and empirical evidence (Zeng et al., 2021) show that a decoupled scope independent from message-passing can improve the expressivity of MPNN by addressing the over-smoothing issue (Li et al., 2018; Kenta Oono, 2020; Shen et al., 2024), and enhances the computational scalability of the models by alleviating the neighbor explosion problem (Chiang et al., 2019; Zeng et al., 2019). Second, as shown in Figure 2, an excessively large visual perception scope can lead to overcrowding of nodes/edges on a fixed-size canvas, resulting in a cluttered appearance that complicates graph interpretation. By using a small $k$, we maintain image clarity independent of message passing.
@@ 5.3 Ablation and Sensitivity Analysis
5.4 Comparison with Positional Encodings
Except comparing with heuristic SFs, to further exhibit the effectiveness of VSFs, we here compared VSFs for nodes with 4 representative encoding mechanisms: 1) 2-d axis position of nodes in subgraph image, 2) Laplacian positional encoding, 3) Distances to other nodes, and 4) Node degree (centrality encoding). These positional encodings (PE) were used as node features and decoded by a 2-layer GCN with a 2-layer MLP for link prediction. The results are presented in Table 10. As can be seen, VSFs outperform the PEs across datasets. Notably, directly encoding the 2-dimensional axis is not effective enough, underscoring the importance of using a vision encoder for comprehensive structural information.
Cora Citeseer Pubmed
VSF 71.32$\pm$0.70 60.96$\pm$0.68 48.27$\pm$0.51
2-D Axis in Image 38.57 $\pm$ 1.22 33.29 $\pm$ 1.80 25.34 $\pm$ 0.95
Laplacian PE 55.14 $\pm$ 0.84 52.03 $\pm$ 1.17 46.80 $\pm$ 0.75
Distance 42.03 $\pm$ 0.82 56.65 $\pm$ 0.43 44.21 $\pm$ 0.36
Degree/Centrality 42.80 $\pm$ 1.52 44.15 $\pm$ 1.61 34.90 $\pm$ 1.23
Table 10: Link prediction effectiveness between VSFs and positional encodings (Hits@100).
@@ 6 Conclusion
Acknowledgement
This work was supported by National Key Research and Development Program of China (No. 2022YFA1004102), NSFC grant (No. 62136005), and the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16200021, 16202523, C7004-22G-1). We also acknowledge the authors of (Chen & Wei, 2025; Lu et al., 2023, 2024, 2025).